# Pavlok-Nudge: A Feedback Mechanism for Atomic Behaviour Modification with Snoring Usecase

### v3

@@ III-C Comparison with Existing BITs
[14]
[17]
1D CNN
8-16 1,337 0.96 0.97 0.98 0.97 0.94
  • O/p chnl: # channels in the output of convolutional layers; P: Precision; R: Recall
IV Discussion and Impact Statement
Impact. While snoring may seem like a minor inconvenience, it can have a significant impact on a person’s health and well-being. One of the most difficult side effects of snoring is sleep disruption, which can lead to daytime fatigue and sleepiness, making it difficult to concentrate and perform daily tasks efficiently. Additionally, snoring can be a sign of sleep apnea, a condition that causes a person to stop breathing periodically throughout the night. This can have serious health implications, including an increased risk of high blood pressure, heart disease and stroke. Snoring can also cause tension and frustration in a sleeping partner, leading to relationship issues and affecting their health by disrupting their sleep. Furthermore, snoring can cause cognitive impairment and memory loss due to a decrease in oxygen levels during sleep. Overall, snoring can have a significant impact on a person’s physical, emotional, and social well-being, and it is important to address the issue promptly to prevent further health complications. Tools such as ours can be useful to assist treatment.
@@ IV Discussion and Impact Statement
Limitation. While the neural network utilised in our proposed system for snoring detection provides accurate and reliable results, one of the limitations is that it is trained on general snoring clips and cannot differentiate between specific users. As a result, in scenarios where someone is sleeping next to their partner, the Pavlok device might be triggered to deliver a stimulus even if the user is not the one snoring. To tackle the problem, we will integrate speaker diarisation [25] to identify the person from whom the audio is coming in a multi-person scenario. The user interface can be revamped to make it more user-friendly. We do not claim that this method can clinically or medically treat snoring. Instead, we present a tool to assist in the personal management of snoring. The tool will need to be clinically tested before we can make any such claims.
@@
Ethical Use and Privacy Concerns. Users are expected to respect the licenses of the deployed components used in Pavlok-Nudge. During run-time, Pavlok-Nudge detects and tracks audio signals in the user’s living space. This could potentially breach the privacy of the subjects present in the environment. Thus, the user must obtain consent from the subjects before using our application. While developing this solution, we make sure that Pavlok-Nudge does not save any audio in the file system. Moreover, all the operations are performed in real-time; buffered audio signals are removed immediately after prediction.
V Conclusion
To the best of our knowledge, the proposed Pavlok-Nudge workflow is one of the first interactive behaviour intervention strategies for atomic behaviour modification. At its core, a deep learning model detects the targeted behavioural pattern and nudges the user through a Pavlok device. Our experiments on the snoring usecase results in commendable performance – achieving SOTA test performance – using a lightweight 1D CNN architecture of only 1,337 trainable parameters. The proposed software application can be easily transferable across different OS platforms. We believe that Pavlok-Nudge have the potential to significantly impact the future behaviour intervention technologies.
@@ Acknowledgement
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and Government of Western Australia. We thank Dr Susannah Soon, Senior Lecturer at Curtin University, for her comments on an initial version of this paper.

### v5

@@ 4.4 Ablation Study
4.5 Limitation
@@
While the neural network utilised in our proposed system for snoring detection provides accurate and reliable results, one of the limitations is that it is trained on general snoring clips and cannot differentiate between specific users. As a result, in scenarios where someone is sleeping next to their partner, the Pavlok device might be triggered to deliver a stimulus even if the user is not the one snoring. To tackle the problem, future work can integrate speaker diarisation [21] to identify the person from whom the audio is coming in a multi-person scenario. The user interface can be revamped to make it more user-friendly. While the current study primarily focuses on the technical validation of the snore detection mechanism, empirical validation in practical settings lies beyond its scope. Future work shall include comprehensive user studies to evaluate the intervention’s effectiveness and to explore user perceptions. We do not claim that this method can clinically or medically treat snoring. Instead, we present a tool that can be clinically tested to assist in the personal management of snoring.
4.6 Ethical Use and Privacy Concerns
@@
Users are expected to respect the licenses of the deployed components used in Pavlok-Nudge. During run-time, Pavlok-Nudge detects and tracks audio signals in the user’s living space. This could potentially breach the privacy of the subjects present in the environment. Thus, the user must obtain consent from the subjects before using our application. While developing this solution, we make sure that Pavlok-Nudge does not save any audio in the file system. Moreover, all the operations are performed in real-time; buffered audio signals are removed immediately after prediction.
5 Conclusion
This paper proposes Pavlok-Nudge, an interactive behaviour intervention technology (BIT) for atomic behaviour modification grounded in aversion therapy. At its core, a deep learning model detects the targeted behavioural pattern, such as snoring, and nudges the user through the Pavlok device. While snoring may seem like a minor inconvenience, it can have a significant impact on a person’s health and well-being, or can be a sign of sleep apnea. One of the most difficult side effects of snoring is sleep disruption, which can lead to daytime fatigue and sleepiness. Our experiments on a public snoring benchmark result in the SOTA test performance using a lightweight 1D CNN architecture of only 1,337 trainable parameters. Apart from snoring, the Pavlok-Nudge framework has the potential to support other targeted behaviour modifications.
@@ Acknowledgement
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia. We thank Dr Susannah Soon, A/Professor at Curtin University, for her comments on an initial version of this paper.